Suppose Fiat recently entered into an Agreement and Plan of Merger with Case for $4.3 billion. Prior to the merger, the market for four-wheel-drive tractors consisted of five firms. The market was highly concentrated, with a Herfindahl-Hirschman index of 3,085. Case’s share of that market was 11 percent, while Fiat comprised just 7 percent of the market. If approved, by how much would the postmerger Herfindahl-Hirschman index increase? Based only on this information, is the Justice Department likely to challenge the merger according to the Horizontal Merger Guidelines?
Answer #1

Original HHI is 3085

New HHI after merger would be calculated as below

We have 5 firms in this concentrated market let's say the share of these other companies be "a";"b";"c" and share for Fiat is 7% , share for Case is 11%

a^2+b^2+c^2+11^2+7^2=3085

a^2+b^2+c^2=2915

When merger between Case and Ford happens then new HHI is a^2+b^2+c^2+18^2=2915+324=3239

Hence the HHI is increases by 3239-3085=154

As this increment in HHI is less than 200 this merger would not likely to be challenged.